# Reliability-Latency-Rate Tradeoff in Low-Latency Communications with Finite-Blocklength Coding

### v3

@@ I Introduction
With the tremendous development of communication technologies, the flow of wireless data transmission has experienced an exponential growth in recent years supported by a higher transmission rate and ubiquitous connectivity. Under this trend, multiple delay-sensitive applications have been gaining significant attention in the fifth-generation (5G) communication networks. Besides, the ongoing works for six-generation (6G) communication networks indicate that there will be more stringent delay and reliability requirements in 6G applications [1]. As one of the essential requirements for mission-critical and emerging applications in 5G and 6G, low-latency communication has been a popular topic under active consideration in the research community.
@@
Among the research for low-latency communications, ultra-reliable and low latency communications (URLLC) is a typical application scenario. URLLC aims to achieve an ultra-low and bounded delay to support the above applications. For the next generation URLLC, the requirement on latency will be more stringent [2]. Additionally, to provide real-time experiences for users, communication latency should be controlled or deterministic [3]. However, due to the fading nature of wireless channels, guaranteeing a bounded delay is usually challenging [4], [5]. Besides, limited maximum power brings more difficulty in achieving a bounded delay [6]. Therefore, statistical quality-of-service (QoS) becomes one of the most common choices as the reliability and latency metrics for URLLC [7]. Statistical QoS is characterized by the queue length or delay thresholds and the corresponding queue length or delay violation probability [8]. As mentioned in [9], focusing solely on average-based metrics without considering the tail behavior of wireless systems makes it difficult to satisfy the low-latency and reliability requirements. There are some works focusing on the characterization of the statistical delay in closed-loop control [10] and edge computing systems [11]. Thus, it is crucial to pay attention to statistical QoS analysis in URLLC and other low-latency systems to characterize the performance limits, which can, in turn, help optimize the system’s QoS performance.
@@
To analyze the QoS of communication systems, the first step is to formulate a proper model, which takes QoS metrics into account. In low-latency communications, we are interested in the probability of the queue length and delay violating certain thresholds [9], which are two key performance indicators of QoS. To this end, the effective bandwidth (EB) and effective capacity (EC) methods are proposed to conduct an asymptotic QoS analysis by introducing the link-layer queueing model. With the help of EB and EC, the queue-length-violation probability (QVP) and delay-violation probability (DVP) can be expressed concisely. The EB of a source is defined as the minimum constant service rate required by the given arrival process to satisfy the QoS requirements [12]. Effective capacity (EC) is the dual concept of EB, indicating the maximum constant source rate that can be supported by the given service process to meet the QoS requirement [13]. Based on EB, EC, and Shannon’s formula, several works focus on the performance analysis and resource allocation designs under statistical QoS constraints in various communication systems with different channel and arrival models, such as independent Rayleigh fading channel [14], time-correlated Rayleigh fading channel [15], and Nakagami-$m$ channel [16]. Specifically, the high-SNR slope of EC was characterized in [14] with the infinite blocklength assumption.
@@
With a proper analysis model, the next step is to formulate the appropriate relationship between rate and reliability in the FBL regime. A commonly quoted requirement for 5G is a latency of less than 1 ms and a packet loss probability that is less than or equal to $10^{-5}$ [7]. Given the limited bandwidth, the assumption that the blocklength shall tend to infinity has to be revised to satisfy these requirements. In particular, the traditional Shannon’s formula with an infinite blocklength is not accurate enough in this regime. Thus, finite-blocklength coding (FBC) is expected to play a central role in addressing this problem. A milestone work [17] presented the maximum coding rate in the FBL regime. The normal approximation was also proposed in [17] to approximate the maximum coding rate for the AWGN channel in the FBL regime. This conclusion inspired further studies on the coding rate of various communication scenarios in the FBL regime, such as Gilbert-Elliott channels [18], quasi-static multiple-antenna fading channels [19], and coherent block fading channels [20]. Besides, Lancho, Koch, and Durisi proposed a high-SNR approximation of the maximum coding rate in the FBL regime based on unitary space-time modulation [21]. With the FBL assumption, EC’s properties differ from those in scenarios with the infinite blocklength assumption.
Combining the FBC analysis with the EC-based model, many works tried to characterize and optimize the performance of low-latency communication systems in terms of reliability, latency, and rate. The authors of [22] validated that the DVP derived from LDT is the upper bound of the actual DVP for certain types of arrival processes, even with a small delay threshold. This result indicated the potential use of the EC-based model in URLLC systems. Based on this conclusion, a joint uplink and downlink resource configuration problem was formulated in [23] to improve the spectrum usage efficiency while ensuring reliability and latency performance. In contrast to these works, another line of research focuses on maximizing EC to improve QoS-constrained throughput in the FBL regime. EC maximization has been studied in various fading channels, including Rayleigh [24, 25], Rician [26], and Nakagami-$m$ [27], each with different constraints and conditions in the FBL regime. In [28], the author determined the throughput of a simple ARQ mechanism in the FBL regime, which was applied in [29]. Although many studies have adopted EC with FBC for QoS optimization in low-latency communication systems, the performance limit of EC in the FBL regime remains an open question. Additionally, due to the channel dispersion term in the FBC rate, deriving the analytical expression of EC is not straightforward. This complexity also poses challenges in evaluating the QoS-constrained throughput of low-latency systems and further optimizing EC.
As mentioned above, although many works focus on optimizing the reliability, latency, and service rate in low-latency systems, there is a lack of comprehensive analysis of the fundamental tradeoff, i.e., reliability-latency-rate tradeoff, in the FBL regime. In our previous work [30], the gain-conservation equation was derived to characterize the relationships between the reliability gain, real-time gain, and service-rate gain. However, in [30], only bounded random arrivals were considered, which is equivalent to the deterministic arrival under the high-SNR assumption. Moreover, the conclusions in [30] did not discuss the influence of high SNR on the FBC. Thus, there is still a need for rigorous analysis of the reliability-latency-rate tradeoff in low-latency communication systems with FBC and random arrivals.
In this paper, we characterize the fundamental tradeoff between reliability, latency, and rate in low-latency systems with FBC. Specifically, we start the analysis from the analysis for the high-SNR slope of EC in the FBL regime. We consider two classical scenarios including transmission over an AWGN channel and a Nakagami-$m$ fading channel. First, we consider the case in which the error probability is fixed to determine the service-rate gain, which is the derivative of the EC to the logarithm of the SNR. Based on this result, we further consider the case in which the error probability is a function of SNR. The reliability gain and real-time gain are then defined as the derivatives of logarithms of error probability and DVP to the logarithm of the SNR. Then, the revised gain-conservation equation is characterized over the AWGN channel and fading channel. Moreover, a discussion of the tradeoff between the error probability and latency is provided for the AWGN channel. For efficiently evaluating the EC in the FBL regime, a Laplace’s-method-based EC approximation approach is proposed to derive the analytical expression of EC. Based on the discussion on the fixed-power allocation scheme, we further propose the EC approximation for certain kinds of channel-gain-based power allocation schemes, while its generalization to the systems with the simple retransmission mechanism is also presented.
Throughout this paper, $u^{\prime}(x)$ denotes the first-order derivative of $u(x)$ for $x$, while $u^{\prime\prime}(x)$ denotes the second-order derivative. $\mathbb{N}^{+}$ denotes the set of positive natural numbers. $\mathbb{E}_{X}\{X\}$ is the expectation of random variable $X$, while ${\rm Var}_{X}\{X\}$ is its variance. For $X[n]$, we omit index $n$ in the subscripts. $e$ denotes Euler’s number. $\boldsymbol{I}_{N}$ denotes a $N$-dimensional identity matrix. For two random variables $X$ and $Y$, ${\rm Cov}(X,Y)$ denote the covariance between $X$ and $Y$. The Gaussian $Q$-function is $Q(x)=\frac{1}{\sqrt{2\pi}}\int_{x}^{\infty}e^{-\frac{t^{2}}{2}}{\rm d}t$, while $Q^{-1}(\cdot)$ represents the inverse Gaussian $Q$-function. $\Gamma(s)=\int_{0}^{\infty}t^{s-1}e^{-t}{\rm d}t$ denote the gamma function. Besides, define the incomplete gamma function as $\Gamma(s,x)=\int_{x}^{\infty}t^{s-1}e^{-t}{\rm d}t$, and the exponential integral function as ${\rm Ei}(x)=-\int_{-x}^{\infty}\frac{e^{-t}}{t}{\rm d}t$.
@@ II System Model and Main Results
Figure 1: System Model.
In this section, we present a system model with FBC. Since we mainly focus on the statistical delay performance of this system, we adopt QVP and DVP as the metrics for evaluating the latency performance. More specifically, we will first describe the system model in Section II-A, which includes physical layer and queueing models. Then, we will introduce the main results of this paper, summarized by four main theorems in Section II-B.
@@ II-A System Model
Next, we introduce the notion of the channel code. In this paper, we assume that the channel state information (CSI) is perfectly known at both the receiver and the transmitter (CSIRT) through the channel estimation and feedback as shown in Fig. 1.222To ensure the perfect CSI assumption, the average power of pilot symbols should be high or the pilot length should be large [32]. With a large coherence bandwidth or coherence time, the pilot length can be large, such as in indoor scenarios with small delay and Doppler spreads [33, 34]. Additionally, to achieve perfect CSIT, there must be a feedback path with a high-resolution quantization scheme and high SNR to transmit CSI precisely and reliably [34]. For latency-constrained systems, available time resources for transmission is limited. Thus, using high SNR or considering systems with a large coherence bandwidth can ensure the perfect CSI assumption while meeting the latency constraints. $\Delta N=B\Delta T$ channel uses are utilized for the channel estimation in each frame, which we neglect in this paper. [^3] Note that for short packet communication with extremely small blocklength, the assumption that perfect CSIRT may be unrealistic. Moreover, analyzing the system with extremely small blocklength requires more effort to derive a precise expression for the coding rate based on [17]. With CSIRT, the transmitter can adaptively determine the coding rate. Specifically, with $h[n]=h$, an $(M_{h},N,\gamma,\epsilon)$ code consists of [35]
@@
    ${\color[rgb]{0,0,0}\Pr\left\{g_{\rm rx}^{h}(\boldsymbol{x},h)\neq J|J\in\{1,\cdots,M_{h}\}\right\}\leq\epsilon.}$ (4)
@@
$R_{N}^{\epsilon}[n]=\sup\left\{\log_{2}M_{h}:\forall(M_{h},N,\gamma,\epsilon)\text{ code}\,|\,h[n]=h\right\}.$ (5)
@@
where $V[n]$ is the channel dispersion, and $O(1)$ is the error term that is bounded for large $N$ and may be related to $\gamma$ and $\epsilon$ at the receiver [17]. [^4] Note that there is no existing work presenting a general conclusion on whether the error term is bounded for $\gamma$. We will discuss this in Section II-B. For the AWGN channel, we omit the index $n$. According to [36], the maximum coding rate (in bits per $N$ channel uses) for a given $\epsilon$, $N$, and $\gamma$ for the AWGN channel is given by
${\color[rgb]{0,0,0}R_{N}^{\epsilon}=N\log_{2}(1+\gamma)-\sqrt{N\left(1-\frac{1}{(1+\gamma)^{2}}\right)}Q^{-1}(\epsilon)\log_{2}e+\frac{\log_{2}N}{2}+G(N,\gamma,\epsilon),}$ (7)
where we call $G(N,\gamma,\epsilon)$ error term function. For the given $\gamma$ and $\epsilon$, $G(N,\gamma,\epsilon)$ satisfies [36]
${\color[rgb]{0,0,0}\lim_{N\to+\infty}|G(N,\gamma,\epsilon)|<+\infty.}$ (8)
Moreover, given the channel gain $|h[n]|^{2}$, the maximum coding rate (in bits per $N$ channel uses) with $\epsilon$, $N$, and $\gamma$ for the fading channel is given by
${\color[rgb]{0,0,0}R^{{}_{\epsilon}}_{N}[n]\!\!=\!\!N\log_{2}\!\left(1\!+\!|h[n]|^{2}\gamma\right)\!-\!\sqrt{\!N\!\left(1\!-\!\frac{1}{\left(1+|h[n]|^{2}\gamma\right)^{2}}\right)}Q^{-1}(\epsilon)\log_{2}e+\frac{\log_{2}N}{2}+G\left(N,|h[n]|^{2}\gamma,\epsilon\right).}$ (9)
For simplicity, we define $\tilde{R}^{\epsilon}_{N}[n]=R^{\epsilon}_{N}[n]-G(N,|h[n]|^{2}\gamma,\epsilon)$, which is called normal approximation in [17].
@@
Remark 1: In this paper, we mainly focus on the analysis with CSIRT. There are also some conclusions on the coding rate in the FBL regime with CSIR only. The rigorous analysis of $R_{N}^{\epsilon}$ in the block fading channel with CSIR for single-input and single-out (SISO) systems was performed in [37]. With real transmitted symbols and noise, an analytical expression of the dispersion was given in Eqs. (27) and (36) of [37] under the assumption that each symbol experiences $i.i.d.$ fading. For simplicity, we omit the index $n$ in this remark. The dispersion for the real block fading channel is given by
${\color[rgb]{0,0,0}V_{r}={\rm Var}_{|h|^{2}}\left(\frac{1}{2}\log_{2}\left(1+|h|^{2}\gamma\right)\right)+\frac{\log_{2}e}{2}\left(1-\mathbb{E}_{|h|^{2}}^{2}\left\{\frac{1}{1+|h|^{2}\gamma}\right\}\right).}$ (11)
@@
${\color[rgb]{0,0,0}V_{c}={\rm Var}_{|h|^{2}}\left(\log_{2}\left(1+|h|^{2}\gamma\right)\right)+\log_{2}e\left(1-\mathbb{E}_{|h|^{2}}^{2}\left\{\frac{1}{1+|h|^{2}\gamma}\right\}\right).}$ (12)
This conclusion is generalized in [20], which considered multiple-input and multiple-output (MIMO) systems with real transmitted symbols and noise. Additionally, they considered the case in which the $T_{d}$ symbols experience the same fading. According to Eqs. (16) and (17) in [20], the dispersion $V$ for a SISO system with a real block fading channel is denoted by
@@
$=$ $T_{d}{\rm Var}_{|h|^{2}}\left(\frac{1}{2}\log_{2}\left(1+|h|^{2}\gamma\right)\right)+\frac{\log_{2}^{2}e}{2}\left(1-\mathbb{E}_{|h|^{2}}^{2}\left\{\frac{1}{1+|h|^{2}\gamma}\right\}\right),$
where $\eta_{1}$, $\eta_{2}$, $v^{*}(1,T_{d})$ are provided in Eqs. (14), (15), and (23) of [20], respectively.
By comparing the results from [37] and [20], we can see that the result in [20] is a generalization of [37] for the real block fading channel. Given $T_{d}=1$, they are consistent. For more general cases, e.g., $T_{d}>1$ in SISO systems or MIMO systems with real block fading channel, the result in [20] provides further insights.

### v4

@@ I Introduction
With the tremendous development of communication technologies, the flow of wireless data transmission has experienced an exponential growth in recent years, supported by a higher transmission rate and ubiquitous connectivity. Under this trend, multiple delay-sensitive applications have garnered significant attention in the fifth-generation (5G) communication networks. Furthermore, ongoing research on the six-generation (6G) communication networks indicates that there will be more stringent requirements for delay and reliability in 6G applications [1]. As one of the essential requirements for mission-critical and emerging applications in 5G and 6G, low-latency communication has been a popular topic under active consideration in the research community.
@@
In the realm of low-latency communications, ultra-reliable and low-latency communications (URLLC) represents a typical application scenario. URLLC aims to achieve an ultra-low and bounded delay to support mission-critical applications. For the next-generation URLLC, the latency requirement will be more stringent [2]. Additionally, providing real-time user experiences requires communication latency to be tightly controlled or deterministic [3]. However, due to the fading nature of wireless channels, ensuring a bounded delay is usually challenging [4], [5]. Furthermore, limited maximum power brings more difficulty of achieving a bounded delay [6]. As a result, statistical quality-of-service (QoS) metrics have become the prevalent choice for assessing reliability and latency in URLLC [7]. Statistical QoS is typically characterized by the queue length or delay thresholds and the associated queue length or delay violation probability [8]. As mentioned in [9], focusing solely on average-based metrics without considering the tail behavior of wireless systems makes it difficult to satisfy the low-latency and reliability requirements. Several studies have explored the statistical delay in closed-loop control [10] and edge computing systems [11]. Therefore, incorporating statistical QoS analysis into URLLC and other low-latency systems is essential for characterizing performance limits and optimizing QoS metrics accordingly.
@@
To analyze the QoS of communication systems, the first step is to formulate a proper model, which takes QoS metrics into account. In low-latency communications, the primary concern lies in the probabilities of queue length and delay exceeding certain thresholds [9], both of which are critical indicators of QoS performance. To this end, the concepts of effective bandwidth (EB) and effective capacity (EC) are utilized to perform asymptotic QoS analysis via a link-layer queuing model. With the help of EB and EC, the queue-length-violation probability (QVP) and delay-violation probability (DVP) can be expressed concisely. EB refers to the minimum constant service rate required for a given arrival process to meet the QoS constraints [12]. EC is the dual concept of EB, representing the maximum constant source rate that can be supported by the given service process to meet the QoS requirement [13]. Based on EB, EC, and Shannon’s formula, several studies focused on the performance analysis and resource allocation designs under statistical QoS constraints in various communication systems with different channel and arrival models, including independent Rayleigh fading channel [14], time-correlated Rayleigh fading channel [15], and Nakagami-$m$ channel [16]. Specifically, the high-SNR slope of EC was analyzed in [14] under the assumption of infinite blocklength.
@@
With a proper analysis model established, the next step is to reveal the relationship between rate and reliability in the finite blocklength (FBL) regime. A commonly quoted requirement for 5G is achieving a latency of less than 1 ms and a packet loss probability no greater than $10^{-5}$ [7]. To satisfy these requirements, the assumption that the blocklength shall tend to infinity has to be revised given the limited bandwidth. In particular, Shannon’s formula, which is asymptotically accurate with the infinite blocklength, is not accurate enough in the FBL regime. Thus, the precise approximation of the capacity with finite-blocklength coding (FBC) is expected to play a central role in addressing this problem. A milestone work [17] analyzed the maximum coding rate in the FBL regime. The normal approximation was also proposed in [17] to estimate the maximum coding rate for the AWGN channel in the FBL regime. This conclusion inspired further studies into the analysis of the coding rate for various communication scenarios in the FBL regime, including Gilbert-Elliott channels [18], quasi-static multiple-antenna fading channels [19], and coherent block fading channels [20]. Additionally, Lancho, Koch, and Durisi proposed a high-SNR approximation of the maximum coding rate in the FBL regime based on unitary space-time modulation [21]. Under the FBL assumption, the properties of EC differ from those derived under the infinite blocklength assumption.
@@
Combining the FBC analysis with the EC-based model, many studies tried to characterize and optimize the performance of low-latency communication systems in terms of reliability, latency, and rate. There are mainly two lines of work on this topic. The first line of work utilized EC to derive the QVP and DVP of the low-latency systems and then optimized the system performance. The authors of [22] validated that the DVP derived from EC serves as the upper bound of the actual DVP for certain types of arrival processes, even with small delay thresholds. This finding highlighted the potential applicability of the EC-based model in URLLC systems. Building on this insight, a joint uplink and downlink resource configuration problem was formulated in [23], aiming to enhance the spectrum usage efficiency while maintaining stringent reliability and latency requirements. In contrast to these works, another line of research focused on maximizing EC to improve QoS-constrained throughput in the FBL regime. EC maximization has been explored in various kinds of fading channels, including Rayleigh [24, 25], Rician [26], and Nakagami-$m$ [27], each considering different constraints and conditions in the FBL regime. In [28], the author calculated the throughput for a simple automatic repeat request (ARQ) mechanism in the FBL regime, which was subsequently applied in [29]. Despite the growing body of research using EC and FBC for QoS optimization in low-latency systems, the theoretical performance limit of EC in the FBL regime remains an open problem. Additionally, the presence of the channel dispersion term in the FBC rate complicates the derivation of an analytical expression for EC, posing additional challenges in assessing QoS-constrained throughput and further refining EC optimization.
As mentioned above, although many works have sought to optimize the reliability, latency, and service rate in low-latency systems, there remains a gap in the comprehensive analysis of the fundamental tradeoff, namely the reliability-latency-rate tradeoff, in the FBL regime. In our previous work [30], the gain-conservation equation was derived to characterize the relationships between the reliability gain, real-time gain, and service-rate gain. However, [30] only considered bounded random arrivals, which is equivalent to the deterministic arrival under the high-SNR assumption. Moreover, the conclusions in [30] did not account for the influence of high SNR on the FBC rate. Therefore, there is still a need for a rigorous analysis of the reliability-latency-rate tradeoff in low-latency communication systems, particularly with FBC and random arrivals.
Figure 1: System Model.
In this paper, we characterize the fundamental tradeoff between reliability, latency, and rate in low-latency systems with FBC. Specifically, we begin by analyzing the high-SNR slope of EC in the FBL regime. We consider two classical scenarios including transmission over an AWGN channel and a Nakagami-$m$ fading channel. First, we consider the case in which the error probability is fixed to determine the service-rate gain, defined as the derivative of EC with respect to the logarithm of the SNR. Based on this result, we further explore the case in which the error probability is a function of SNR. The reliability gain and real-time gain are then defined as the derivatives of logarithms of error probability and DVP with respect to the logarithm of the SNR. Subsequently, we derive the revised gain-conservation equations for both AWGN and fading channels. Additionally, a detailed discussion of the tradeoff between the error probability and latency is provided for the AWGN channel. For efficiently evaluating EC in the FBL regime, we propose a Laplace’s-method-based EC approximation approach to derive an analytical expression for EC. Based on the discussion on the fixed-power allocation scheme, we further extend our analysis to consider the EC approximation for certain kinds of channel-gain-based power allocation schemes, with a generalization to systems incorporating the simple retransmission mechanism.
Throughout this paper, $u^{\prime}(x)$ denotes the first-order derivative of $u(x)$ for $x$, while $u^{\prime\prime}(x)$ denotes the second-order derivative. $\mathbb{N}^{+}$ denotes the set of positive natural numbers. $\mathbb{E}_{X}\{X\}$ is the expectation of random variable $X$, while ${\rm Var}_{X}\{X\}$ is its variance. For $X[n]$, we omit index $n$ in the subscripts. $e$ denotes Euler’s number. $\boldsymbol{I}_{N}$ denotes a $N$-dimensional identity matrix. For two random variables $X$ and $Y$, ${\rm Cov}(X,Y)$ denote the covariance between $X$ and $Y$. The Gaussian $Q$-function is $Q(x)=\frac{1}{\sqrt{2\pi}}\int_{x}^{\infty}e^{-\frac{t^{2}}{2}}{\rm d}t$, while $Q^{-1}(\cdot)$ represents the inverse Gaussian $Q$-function. $\Gamma(s)=\int_{0}^{\infty}t^{s-1}e^{-t}{\rm d}t$ denote the gamma function. Besides, define the incomplete gamma function as $\Gamma(s,x)=\int_{x}^{\infty}t^{s-1}e^{-t}{\rm d}t$, and the exponential integral function as ${\rm Ei}(x)=-\int_{-x}^{\infty}\frac{e^{-t}}{t}{\rm d}t$. $W(\cdot)$ denotes the Lambert W function [31]. $O(\cdot)$ is the big-O notation. For two functions $f(x)$ and $g(x)$, $f(x)=O(g(x))$ holds if there exists a positive real number $M$ and a real number $x_{0}$ such that $|f(x)|\leq M|g(x)|$ for all $x\geq x_{0}$.
@@ II System Model and Main Results
In this section, we present a system model with FBC, focusing on the statistical delay performance. To evaluate the latency performance, we adopt two key metrics: the queue-length-violation probability (QVP) and delay-violation probability (DVP). Specifically, we will first describe the system model in Section II-A, which includes physical layer and queueing models. Following that, we will introduce the main results of this paper, summarized in four main theorems in Section II-B.
@@ II-A System Model
$R_{N}^{\epsilon}=N\log_{2}(1+\gamma)-\sqrt{N\left(1-\frac{1}{(1+\gamma)^{2}}\right)}Q^{-1}(\epsilon)\log_{2}e+\frac{\log_{2}N}{2}+G(N,\gamma,\epsilon).$ (7)
$R^{\epsilon}_{N}[n]=N\log_{2}\left(1+|h[n]|^{2}\gamma\right)-\sqrt{N\left(1-\frac{1}{\left(1+|h[n]|^{2}\gamma\right)^{2}}\right)}Q^{-1}(\epsilon)\log_{2}e+\frac{\log_{2}N}{2}+G\left(N,|h[n]|^{2}\gamma,\epsilon\right).$ (9)
Next, we introduce the notion of the channel code. In this paper, we assume that the channel state information (CSI) is perfectly known at both the receiver and the transmitter (CSIRT) through the channel estimation and feedback as shown in Fig. 1.222To ensure the perfect CSI assumption, either the average power of pilot symbols should be high, or the pilot length should be large [33]. In environments with large coherence bandwidth or coherence time, such as indoor scenarios with small delay and Doppler spread, a large pilot length is feasible [34, 35]. Additionally, achieving perfect CSIT requires a feedback path with high-resolution quantization and sufficient SNR to transmit the CSI accurately and reliably [35]. In latency-constrained systems, available time resources for transmission is limited. Therefore, using high SNR or considering systems with a large coherence bandwidth can help maintain the perfect CSI assumption while adhering to latency constraints. $\Delta N=B\Delta T$ channel uses are utilized for the channel estimation in each frame, which we neglect in this paper. [^3] Note that for short packet communication with extremely small blocklength, the assumption that perfect CSIRT may be unrealistic. Moreover, analyzing the system with extremely small blocklength requires more effort to derive a precise expression for the coding rate based on [17]. With CSIRT, the transmitter can adaptively determine the coding rate. Specifically, with $h[n]=h$, an $(M_{h},N,\gamma,\epsilon)$ code consists of [19]
@@
    $\sum_{j=1}^{M_{h}}\Pr\left\{g_{\rm rx}^{h}(\boldsymbol{x},h)\neq J|J=j\right\}\Pr\{J=j\}\leq\epsilon.$ (4)
@@
$R_{N}^{\epsilon}[n]=\sup$$\left\{\log_{2}M_{h}:\right.$ (5)
$\left.\forall(M_{h},N,\gamma,\epsilon)\text{ code}\,\,|\,\,h[n]=h\right\}.$
@@
where $V[n]$ is the channel dispersion, and $O(1)$ is the error term that is bounded for large $N$ and may be related to $\gamma$ and $\epsilon$ at the receiver [17]. [^4] Note that there is no existing work presenting a general conclusion on whether the error term is bounded for $\gamma$. We will discuss this in Section II-B. For the AWGN channel, we omit the index $n$. According to [36], the maximum coding rate (in bits per $N$ channel uses) for a given $\epsilon$, $N$, and $\gamma$ for the AWGN channel is presented in Eq. (7), which is located at the bottom of this page. In Eq. (7), we call $G(N,\gamma,\epsilon)$ error term function. For the given $\gamma$ and $\epsilon$, $G(N,\gamma,\epsilon)$ satisfies [36]
$\lim_{N\to+\infty}|G(N,\gamma,\epsilon)|<+\infty.$ (8)
Moreover, given the channel gain $|h[n]|^{2}$, the maximum coding rate (in bits per $N$ channel uses) with $\epsilon$, $N$, and $\gamma$ for the fading channel is given in Eq. (9), which is presented at the bottom of this page. For simplicity, we define $\tilde{R}^{\epsilon}_{N}[n]=R^{\epsilon}_{N}[n]-G(N,|h[n]|^{2}\gamma,\epsilon)$, which is called normal approximation in [17].
@@
Remark 1: In this paper, we primarily focus on the analysis with CSIRT. There are also some existing conclusions on the coding rate in the FBL regime with CSIR only. A rigorous analysis of $R_{N}^{\epsilon}$ in the block fading channel with CSIR for single-input and single-out (SISO) systems was carried out in [37]. For real transmitted symbols and noise, an analytical expression for the dispersion was given in Eqs. (27) and (36) of [37] under the assumption that each symbol experiences $i.i.d.$ fading. For simplicity, we omit the index $n$ in this remark. The dispersion for the real block fading channel is given by
$V_{r}={\rm Var}_{|h|^{2}}$ $\left(\frac{1}{2}\log_{2}\left(1+|h|^{2}\gamma\right)\right)+$ (11)
$\frac{\log_{2}e}{2}\left(1-\mathbb{E}_{|h|^{2}}^{2}\left\{\frac{1}{1+|h|^{2}\gamma}\right\}\right).$
@@
$V_{c}={\rm Var}_{|h|^{2}}$ $\left(\log_{2}\left(1+|h|^{2}\gamma\right)\right)+$ (12)
$\log_{2}e\left(1-\mathbb{E}_{|h|^{2}}^{2}\left\{\frac{1}{1+|h|^{2}\gamma}\right\}\right).$
This conclusion was further generalized in [20], which considered multiple-input and multiple-output (MIMO) systems with real transmitted symbols and noise. Additionally, they considered the case in which $T_{d}$ symbols experience the same fading. According to Eqs. (16) and (17) in [20], the dispersion $V$ for a SISO system with a real block fading channel is represented by Eq. (13), which is presented at the bottom of this page. In Eq. (13), $\eta_{1}$, $\eta_{2}$, $v^{*}(1,T_{d})$ are defined by Eqs. (14), (15), and (23) of [20], respectively.
@@
$=$ $T_{d}{\rm Var}_{|h|^{2}}\left(\frac{1}{2}\log_{2}\left(1+|h|^{2}\gamma\right)\right)+\frac{\log_{2}^{2}e}{2}\left(1-\mathbb{E}_{|h|^{2}}^{2}\left\{\frac{1}{1+|h|^{2}\gamma}\right\}\right).$
A comparison between the results of [37] and [20] reveals that the findings in [20] serve as a generalization of [37] for the real block fading channel. When $T_{d}=1$, the results align with each other. However, for more general cases, such as $T_{d}>1$ in SISO systems or MIMO systems with real block fading channel, [20] offers additional insights.